OpenStudy (anonymous):

If a^2-64 equals the area of a rectangle. What is the perimeter of the rectangle?

OpenStudy (solomonzelman):

And back to un-exact questions with infinite number of possible answers

OpenStudy (solomonzelman):

You can technically divide `a²-64` not just into (x-8) and (x+8) which is what they are expecting you to do, But, also you can divide it into many other things (although there will be remainders.

OpenStudy (solomonzelman):

Let's just go with `a² - 64` = `(a - 8)` × `(a+8)` ↓ ↓ ↓ `AREA` = `Width` × `LENGTH`

OpenStudy (anonymous):

Ok, now that I have the side lenghts , what do I have to do to get the penrimeter?

OpenStudy (solomonzelman):

Perimeter = (width + Length) × 2

OpenStudy (solomonzelman):

So, now in your case, Perimeter = ( `width` + `length` ) × 2 ↓ ↓ ↓ Perimeter = ( `[x+8]` + `[x-8]` ) × 2

OpenStudy (anonymous):

Would it be 2 a?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ ✂ }\)

OpenStudy (anonymous):

Again, thank you ! You really help me understand how to solve the question.

OpenStudy (solomonzelman):

Good ! Other users just try to get answers on here, you are not like that... I am glads to help you (when I am able to) ! \(\LARGE\color{red }{\bf {Also} }\) \(\normalsize\color{blue}{\bf {you~~~can~~~find~~~any~~~section~~~by:} }\) 1) clicking "more subjects" (the top left corner, towards the center) 2) typing in the name of the section (name of a section can be "history" Mathematics and other section for each unique area of studying). \(\normalsize\color{blue}{\bf {You~~~can~~~FAN~~~someone~~~by:} }\) 1) clicking on the user's name 2) clicking "fan" on the bottom right corner of user's box. ( you can also write a testimonial by clicking "add testimonial" and typing something in testimonial box) \(\normalsize\color{blue}{\bf {You~~~can~~~mention~~~someone~~~in~~~your~~~question} }\) (or in any other question) by writing @ in front of their name, just like I would mention you... LOOK ! .... @LOVEVAL (they would get a notification that a user mentioned them in a reply) \(\normalsize\color{blue}{\bf {You~~~can~~~ message~~~someone~~~by:} }\) 1) clicking on the someone's username 2) clicking "Send Message" 3) type your message press submit, and wait till the messages submits. \(\normalsize\color{blue}{\bf {You~~~can~~~ see~~~if~~~a~~~user~~~is~~~offline~~~by} }\) Clicking on the user's username, and below his picture you will see (in light blue) “offline”, “online” or you might see that he is viewing or replying to a specific question. You might as well see that he is just looking around in a specific section. If you have any questions you can message me or anyone else on this website. We will be happy to respond !